buttercup888:

Please can someone help me The sets of numbers 6, 8, 10 and 5, 12, 13 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples! (5 points)

buttercup888:

@Aeon @AngeI @Bob

AngeI:

Bob is offline and aeon is a ai but I will try to help

buttercup888:

Thank you

AngeI:

I honestly am too tired to explain this im sorry so I will just copy and paste what will help you.

AngeI:

Just write something like this: Let's plug in 6, 8 and 10 into the pythagorean theorem: a^2 + b^2 = c^2 6^2 + 8^2 = 10^2 36 + 64 = 100 100 = 100 So 6, 8, and 10 are pythagorean triples because they satisfy the pythagorean theorem. Let's plug in 5, 12 and 13 into the pythagorean theorem: a^2 + b^2 = c^2 5^2 + 12^2 = 13^2 25 + 144 = 169 169 = 169 So 5, 12, and 13 are also pythagorean triples because they satisfy the pythagorean theorem.

buttercup888:

Thank you

buttercup888:

Can you help me with another one if you dont mind?

AngeI:

I can see if i can google the answer for you bu I am too tired to do anything else

buttercup888:

okay them nvm

buttercup888:

then*

AngeI:

Sorry, but you can most likely find the answer if you try googling it